# LORE: Lagrangian-Optimized Robust Embeddings for Visual Encoders

### v1

@@ Appendix B The Parametrization Gap
Instead of the original constrained optimization problem. in this section we are interested in deriving some bounds on the duality gap, following the proof from Robey et al. [2021].
@@
$d^{*}-\sup_{\lambda\in\Lambda}\inf_{\theta\in\Theta}L(\phi_{\theta},\lambda)\geq\inf_{\theta\in\Theta}L(\phi_{\theta},\tilde{\lambda}^{*})=\inf_{\phi\in\mathcal{H}}L(\phi,\tilde{\lambda}^{*})\geq\inf_{g\in\text{conv}(\mathcal{H})}L(g,\tilde{\lambda}^{*})$
@@
Finally, we note that although we have treated the dual space as the full infinite-dimensional set $\Lambda=\{\lambda:\mathcal{X}\to[0,\infty)\}$, in this work we have restricted $\lambda$ to a finite-dimensional family $\{\lambda_{\omega}\}_{\omega\in\Omega}\subset\Lambda$ that uniformly approximates its elements. Concretely, if for every $\lambda\in\Lambda$ there exists $\omega\in\Omega$ with $\|\lambda-\lambda_{\omega}\|_{L^{1}(\mathcal{D})}\leq\xi$, then replacing
@@
only incurs an arbitrarily small error $O(\xi)$. All weak-duality arguments carry over immediately, and under Slater’s condition the resulting strong-duality statement remains valid up to this negligible approximation.
@@
otherwise $L(\phi,\omega)\to+\infty$ as $\omega\to+\infty$. In other words, the constant-$\lambda$ relaxation exactly enforces
@@
Thus, by limiting the expressivity of $\lambda$, we move from the original per-sample constraint
@@ Appendix E Generalization Gap in Adversarial Training
where =$\mathcal{L_{H}}=\{\ell(\phi,\cdot)\mid\phi\in\mathcal{H}\}$ is the loss class induced by hypothesis class $\mathcal{H}$, $\mathfrak{R}_{n}(\mathcal{L_{H}})$ is the empirical Rademacher complexity of $\mathcal{L_{H}}$, and $\sup_{\phi,x}|\ell(\phi,x)|\leq B$.
@@ Appendix F Deviation Between Cosine Similarities
$\frac{u}{\|u\|}-\frac{\hat{u}}{\|\hat{u}\|}=\Bigl{(}\frac{u}{\|u\|}-\frac{\hat{u}}{\|u\|}\Bigr{)}+\Bigl{(}\frac{\hat{u}}{\|u\|}-\frac{\hat{u}}{\|\hat{u}\|}\Bigr{)},$
@@
$\Bigl{\|}\tfrac{u}{\|u\|}-\tfrac{\hat{u}}{\|\hat{u}\|}\Bigr{\|}\;\leq\;\frac{\|u-\hat{u}\|}{\|u\|}+\|\hat{u}\|\;\Bigl{|}\tfrac{1}{\|u\|}-\tfrac{1}{\|\hat{u}\|}\Bigr{|}.$
@@ Appendix H Revisiting Embedding Models
CLIP [Radford et al., 2021].  A major part of our experiments builds upon CLIP, which consists of an image encoder $\phi$ and a text encoder $\psi$ that map images and text descriptions into a shared embedding space. For zero-shot classification, textual descriptions are typically formatted as "This is a photo of a [CLS]”, where [CLS] represents class labels. The probability of assigning an image $x$ to a class $\hat{y}$ is computed via a softmax over cosine similarities:
@@
DINOv2 [Oquab et al., 2024].  In addition to CLIP, we incorporate DINOv2, a powerful self-supervised visual transformer-based encoder, into our exploration of embedding models. While CLIP provides a joint image-text embedding space, DINOv2 focuses solely on visual representation learning. This complementary perspective allows us to compare and leverage both multimodal and unimodal embedding paradigms.
@@ Appendix I Impact of $K$ on Model Performance
In this section, we evaluate the robustness of our fine-tuning approach against a black-box adversarial attack known as the Square Attack [Andriushchenko et al., 2020]. Unlike gradient-based methods, Square Attack operates without access to model gradients and perturbs the input using a query-efficient, score-based strategy. This makes it a strong candidate for evaluating real-world robustness where white-box access is not feasible. We conduct experiments on LORE4, which is adversarially fine-tuned on ImageNet, to assess how well the model generalizes to such black-box settings. The results, summarized in Table 7, show that our method consistently outperforms the baseline under this challenging threat model.
@@
Table 12: Training time (in seconds) for 50 iterations of LORE using 4×H100 GPUs across different architectures for $\lambda_{\omega}$. The Linear model uses a single-layer projection: nn.Linear(self.embedding_size, 1). Lower is better.

### v2

@@ Appendix B The Parametrization Gap
Instead of the original constrained optimization problem. in this section we are interested in deriving some bounds on the duality gap, following the proof from robey2021adversarialrobustnesssemiinfiniteconstrained.
@@
[…]
@@
Finally, we note that although we have treated the dual space as the full infinite‐dimensional set $\Lambda=\{\lambda:\mathcal{X}\to[0,\infty)\}$, in this work we have restricted $\lambda$ to a finite‐dimensional family $\{\lambda_{\omega}\}_{\omega\in\Omega}\subset\Lambda$ that uniformly approximates its elements. Concretely, if for every $\lambda\in\Lambda$ there exists $\omega\in\Omega$ with $\|\lambda-\lambda_{\omega}\|_{L^{1}(\mathcal{D})}\leq\xi$, then replacing
@@
only incurs an arbitrarily small error $O(\xi)$. All weak‐duality arguments carry over immediately, and under Slater’s condition the resulting strong‐duality statement remains valid up to this negligible approximation.
@@
Thus, by limiting the expressivity of $\lambda$, we move from the original per‐sample constraint
@@ Appendix E Generalization Gap in Adversarial Training
where $\mathcal{L_{H}}=\{\ell(\phi,\cdot)\mid\phi\in\mathcal{H}\}$ is the loss class induced by hypothesis class $\mathcal{H}$, $\mathfrak{R}_{n}(\mathcal{L_{H}})$ is the empirical Rademacher complexity of $\mathcal{L_{H}}$, and $\sup_{\phi,x}|\ell(\phi,x)|\leq B$.
@@ Appendix F Deviation Between Cosine Similarities
[…]
@@
[…]
@@ Appendix H Revisiting Embedding Models
CLIP [radford_learning_2021].  A major part of our experiments builds upon CLIP, which consists of an image encoder $\phi$ and a text encoder $\psi$ that map images and text descriptions into a shared embedding space. For zero-shot classification, textual descriptions are typically formatted as "This is a photo of a [CLS]”, where [CLS] represents class labels. The probability of assigning an image $x$ to a class $\hat{y}$ is computed via a softmax over cosine similarities:
@@
DINOv2 [oquab2024dinov2learningrobustvisual].  In addition to CLIP, we incorporate DINOv2, a powerful self-supervised visual transformer-based encoder, into our exploration of embedding models. While CLIP provides a joint image-text embedding space, DINOv2 focuses solely on visual representation learning. This complementary perspective allows us to compare and leverage both multimodal and unimodal embedding paradigms.
@@ Appendix I Impact of $K$ on Model Performance
In this section, we evaluate the robustness of our fine-tuning approach against a black-box adversarial attack known as the Square Attack [andriushchenko2020squareattackqueryefficientblackbox]. Unlike gradient-based methods, Square Attack operates without access to model gradients and perturbs the input using a query-efficient, score-based strategy. This makes it a strong candidate for evaluating real-world robustness where white-box access is not feasible. We conduct experiments on LORE4, which is adversarially fine-tuned on ImageNet, to assess how well the model generalizes to such black-box settings. The results, summarized in Table 9, show that our method consistently outperforms the baseline under this challenging threat model.
@@
J.6 Ablation on Sample-Specific Margin $m(x)$
The choice of a sample-specific tolerance margin, $m(x)=\rho\|\phi_{\theta_{0}}(x)\|^{2}$, is a crucial design decision in the LORE framework.
Using a sample-invariant tolerance ($\rho_{\text{fixed}}$), where the constraint is fixed across all data points ($d(\phi_{\theta}(x),\phi_{\theta_{0}}(x))\leq\rho_{\text{fixed}}$), offers an alternative but is challenging. It requires extensive, task-dependent and model-dependent hyperparameter tuning that severely limits its practical utility and generalization ability.
By contrast, the sample-specific margin provides an adaptive, scale-invariant tolerance tied to the pre-trained embedding’s norm. This design choice reduces manual tuning, preserves semantic structure by relating the constraint geometrically to the cosine similarity of embeddings, and ultimately offers greater stability.
Experimental Comparison. To validate this choice, we conducted experiments using ViT-B/32 CLIP models fine-tuned for 5 epochs on ImageNet-100. We compare the adaptive margin (LOREm(x)) against several empirically chosen fixed tolerances ($\rho=0.6,0.8,1.5$). The superscript ($\varepsilon$) denotes the training perturbation budget.
As shown in Table 13, fixed tolerances of $0.6$ and $1.5$ lead to suboptimal performance, demonstrating that a fixed margin is either overly restrictive (degrading robustness) or overly loose (causing behavior similar to FARE). The sample-specific margin (LOREm(x)) consistently achieves the best overall Average Robust Accuracy, validating its superior stability and efficacy across different training settings.
Table 13: Comparison of Adaptive Margin $m(x)$ vs. Fixed Margins (ViT-B/32 on ImageNet-100). The Avg Robust Acc column is the average of the three adversarial perturbation levels ($\varepsilon=1,2,4$).
Train $\varepsilon$ Method $\rho$ Clean $\varepsilon=1$ $\varepsilon=2$ $\varepsilon=4$ Avg Robust Acc
2 LORE2m(x) 0.1 68.81 51.51 33.71 9.93 31.72
LORE2 0.6 70.12 24.32 18.65 0.56 14.51
LORE2 0.8 69.65 45.65 29.45 5.65 26.92
LORE2 1.5 63.45 41.23 25.47 3.64 23.45
FARE2 62.23 40.12 23.15 3.40 22.22
4 LORE4m(x) 0.1 66.91 46.99 35.71 17.58 33.43
LORE4 0.6 67.21 19.50 7.21 1.23 9.31
LORE4 0.8 64.32 40.21 32.23 14.68 29.04
LORE4 1.5 55.24 33.21 29.32 11.74 24.76
FARE4 53.68 34.85 28.78 12.87 25.50
J.7 LORE in Supervised Adversarial Fine-Tuning
Given an image-label pair $(x,y)$, TeCoA [mao2023understandingzeroshotadversarialrobustness] optimizes a contrastive loss to align adversarial image embeddings with text embeddings. The loss function is defined as:
$L_{\text{sup}}(x,y;\theta)=\max_{\delta\in\Delta}-\sum_{i}\Bigg[y_{i}\log\frac{\exp(\cos(z,\psi(t_{i}))/\tau)}{\sum_{k}\exp(\cos(z,\psi(t_{k}))/\tau)}\Bigg],$ (25)
where $z=\phi_{\theta}(x+\delta)$ is the adversarial image embedding and $\psi(t_{i})$ is the text embedding. Here, $\psi(t_{i})$ corresponds to the text embedding of the $i$-th class, where each class label is inserted into a fixed natural language template (e.g., "This is a photo of {class}").
We incorporated our proximity-based regularizer into TeCoA, forming a combined method ($\text{TeCoA}+l_{2}$), where $l_{2}$ denotes the $\ell_{2}$-based embedding proximity constraint:
$\max_{\omega\in\Omega}\min_{\theta\in\Theta}E_{x,y\sim D}\left[L_{\text{sup}}(x,y;\theta)+\lambda_{\omega}(x)\big(|\phi_{\theta}(x)-\phi_{0}(x)|_{2}^{2}-\rho|\phi_{0}(x)|_{2}^{2}\big)\right]$
Distribution-Level Constraint. We also explored a probability-distribution variant of LORE for the supervised setting, constraining the KL divergence between the fine-tuned model’s output distributions and those of the pretrained model. This approach requires only class text embeddings, not sample labels. Applied to TeCoA, it yields a second regularized variant: $\text{TeCoA}+\text{KL}$.
For classification using probability distributions induced by the contrastive model, we formulate the constrained optimization problem:
$\min_{\theta\in\Theta}E_{x,y\sim D}\big[L_{\text{sup}}(x,y;\theta)\big]\quad\text{s.t.}\quad D_{\text{KL}}(p_{\theta_{0}}(\cdot\mid x),p_{\theta}(\cdot\mid x))\leq\rho\cdot m(x),\quad\text{where}\quad m(x)=H(p_{\theta_{0}}(\cdot\mid x))$
Building on this formulation, LORE’s final objective function can be computed as before. The framework adjusts the proximity constraint by model confidence: low entropy (high confidence) tightens the constraint, while high entropy relaxes it for greater adaptation flexibility.
Results. The experimental setting involved fine-tuning ViT-B/32 CLIP models for 15 epochs on ImageNet-100 using a learning rate of 1e51e-5. The proximity constraint hyperparameter used $\rho=0.1$. The results are shown in Table 14.
TeCoA consistently benefits from both LORE-based ($l_{2}$) and KL regularization, achieving improved robustness while maintaining clean performance. Notably, for high adversarial training budgets (e.g., Train $\varepsilon=8$), the $\text{TeCoA}+l_{2}$ method yields a massive increase in clean accuracy compared to the baseline TeCoA. These results demonstrate that proximity-based constraints—whether applied in embedding space ($l_{2}$) or probability space (KL)—can effectively stabilize adversarial fine-tuning in supervised frameworks as well.
Table 14: Extension to Supervised Fine-Tuning (ViT-B/32 CLIP on ImageNet-100). The Avg Robust Acc column is the average of the four adversarial perturbation levels ($\varepsilon=1,2,4,8$).
Train $\varepsilon$ Method Clean $\varepsilon=1$ $\varepsilon=2$ $\varepsilon=4$ $\varepsilon=8$ Avg Robust Acc
2 TeCoA2 60.04 (ref) 49.11 35.94 14.73 0.50 25.07 (ref)
TeCoA2 + $l_{2}$ 75.97 (+15.93 $\uparrow$) 53.24 39.12 18.12 1.23 27.93 (+2.86 $\uparrow$)
TeCoA2 + KL 66.50 (+6.46 $\uparrow$) 53.04 39.53 16.94 0.60 27.53 (+2.46 $\uparrow$)
4 TeCoA4 49.55 (ref) 44.20 38.21 19.80 1.03 25.81 (ref)
TeCoA4 + $l_{2}$ 73.12 (+23.57 $\uparrow$) 56.20 53.12 49.12 5.23 40.92 (+15.11 $\uparrow$)
TeCoA4 + KL 56.01 (+6.46 $\uparrow$) 48.62 42.03 22.18 1.25 28.52 (+2.71 $\uparrow$)
8 TeCoA8 30.22 (ref) 24.72 19.09 7.41 2.35 13.39 (ref)
TeCoA8 + $l_{2}$ 67.23 (+37.01 $\uparrow$) 36.21 26.40 15.12 4.23 20.49 (+7.10 $\uparrow$)
TeCoA8 + KL 38.06 (+7.84 $\uparrow$) 27.69 21.94 8.75 2.82 15.30 (+1.91 $\uparrow$)
J.8 Validation on Downstream Tasks (Segmentation)
To confirm LORE’s generalizability beyond classification, we evaluated its effectiveness on semantic segmentation, a critical VLM downstream task. We use the open-source benchmark from kowalczuk2024benchmarkingrobustselfsupervisedlearning to have a fair comparison with our baselines.
We assessed the robustness of the DINOv2 ViT-S/14 model on the ADE20k segmentation dataset. The results in Table 15 show that the LORE-hardened model improves robustness against adversarial attacks in the embedding space while maintaining a favorable clean accuracy trade-off.
Table 15: Segmentation Results on ADE20k (DINOv2 ViT-S/14). $\text{LORE}^{8}$ and $\text{FARE}^{8}$ denote models trained with an $\varepsilon=8/255$ perturbation budget. mIoU is reported for clean data and under an adversarial attack in the embedding space.
Method Clean mIoU Embed Attack mIoU
Pretrained 0.42 0.23
$\text{LORE}^{8}$ 0.43 0.35
$\text{FARE}^{8}$ 0.23 0.15
@@
Table 17: Training time (in seconds) for 50 iterations of LORE using 4×H100 GPUs across different architectures for $\lambda_{\omega}$. The Linear model uses a single-layer projection: $n$.Linear(self.embedding_size, 1). Lower is better.